OpenStudy (anonymous):

A particle P moves in a straight line,starting from rest at the point O. At time t seconds after leaving O, the acceleration a ms^-2, of P is given by a=4+12t Calculate the distance traveled by P in the third second.

OpenStudy (anonymous):

Thanks for having a look anyway...

OpenStudy (anonymous):

im lost D: but ill tag sum ppls c: @AccessDenied @anonymous_user @ememlove @ganeshie8

OpenStudy (anonymous):

I know the answer, if that helps... its 48m

ganeshie8 (ganeshie8):

familiar with integrals/derivatives, right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\( a = 4+12t\) \( \implies v(t) - v(0) = \int \limits_0^t 4+12t ~dt\) Since the particle started from rest, \(v(0) = 0\) \( \implies v(t) -0 = \int \limits_0^t 4+12t ~dt\)

ganeshie8 (ganeshie8):

evaluate the integral, what do u get ?

OpenStudy (anonymous):

\[4t+6t ^{2}\]

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

\( v(t) = 4t+6t^2\) \(\implies s(t) -s(0) = \int \limits_0^t 4t+6t^2~dt\) since the particle started from Origin, \(s(0) = 0\) \(\implies s(t) -0 = \int \limits_0^t 4t+6t^2~dt\)

ganeshie8 (ganeshie8):

evaluate the integral

OpenStudy (anonymous):

\[2t ^{2}+3t ^{3}\]

ganeshie8 (ganeshie8):

yup ! thats the displacement function : \(s(t) = 2t^2 + 3t^3\)

OpenStudy (anonymous):

So what needs to be done now...

ganeshie8 (ganeshie8):

Calculate the distance traveled by P in the \(third\) second.

ganeshie8 (ganeshie8):

displaccement function : \(s(t) = 2t^2 + 3t^3\) plugin \(t = 3\), distance travelled in 3rd second : \(s(3) = ?\)

OpenStudy (anonymous):

99

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

But in my book it says 48

OpenStudy (anonymous):

So is this a typo in the book, or is our answer incorrect...

ganeshie8 (ganeshie8):

it would be unwise to call the textbook has a typo, lets check our work again

ganeshie8 (ganeshie8):

btw, point O is at the center (0,0) right ?

OpenStudy (anonymous):

I would assume

OpenStudy (anonymous):

yes it is.

ganeshie8 (ganeshie8):

Found the mistake !

ganeshie8 (ganeshie8):

\(\implies s(t) -0 = \int \limits_0^t 4t+6t^2~dt\)

ganeshie8 (ganeshie8):

^evaluate the integral, what do u get ?

OpenStudy (anonymous):

\[2t ^{2}+3t ^{3}\]

OpenStudy (anonymous):

+c

OpenStudy (anonymous):

maybe?

OpenStudy (anonymous):

I have no idea.....

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

\(\implies s(t) -0 = \int \limits_0^t 4t+6t^2~dt \) \(\implies s(t)= 2t^2 + 2t^3\) right ?

ganeshie8 (ganeshie8):

thats our first mistake, we made another mistake in interpreting what the question is asking us to find out

ganeshie8 (ganeshie8):

Calculate the distance traveled by P in the third second. ^^^^^^^^^^^^^^^^^

ganeshie8 (ganeshie8):

quesiton is about finding the distance travelled between t = 2 and t = 3 only

OpenStudy (anonymous):

t=2?

ganeshie8 (ganeshie8):

so you need to find out : \(\large s(3) - s(2)\) if that makes any sense...

OpenStudy (anonymous):

How did you derive t=2 from the question?

OpenStudy (anonymous):

wouldnt it be t=0 and t=3

ganeshie8 (ganeshie8):

t = 0 and t= 3 wud be the distance travelled "in 3 seconds"

ganeshie8 (ganeshie8):

but the question is asking u to find distance travelled "in the 3rd second"

ganeshie8 (ganeshie8):

see the difference ?

OpenStudy (anonymous):

now i get it...

ganeshie8 (ganeshie8):